# Peierls substitution and Hall motion in exotic Carroll dynamics

### v1

@@ V.2 Symmetries of exotic particles
  • Lifting (V.1) to the evolution space of an exotic Galilean particle,
@@
    the Souriau form (III.5) changes as for DJT in (V.3) [with $\sigma_{DJT}\leadsto\sigma_{exoG}$], leading to :
@@
    A nonzero magnetic field $B\neq 0$ breaks Galilean boost symmetry of an exotic Galilean particle (as seen also directly from in Prop.III.1: the rhs is invariant w.r.t. (V.1)).
@@
$\left\{\begin{array}[]{llccll}\text{for}&\beta_{1}=0&d\Big{(}\hbox{$\textstyle\frac{1}{2}$}eEx_{2}^{2}\Big{)}\beta_{2}&\text{closed}&{\quad\Rightarrow}&\text{symmetry}\\[8.0pt] \text{for}&\beta_{2}=0&eEx_{1}dx_{2}\,\beta_{1}&\text{not closed}&{\quad\Rightarrow}&{\text{broken}}\end{array}\right.\qquad.$ (V.16)
@@
Boosting the exotic Carroll particle along the electric field ${\mathbf{E}}=(0,E)$ generates the conserved quantity
${\mathcal{K}}_{||}=(x_{2}-\theta\,p_{1})+\hbox{$\textstyle\frac{1}{2}$}eE\,x_{2}^{2}=\widetilde{Q}_{2}+\hbox{$\textstyle\frac{1}{2}$}eE\,x_{2}^{2}\,\,.$ (V.17)
The conservation of ${\mathcal{K}}_{||}$ is readily confirmed by using (after replacing $eB$ by $B^{*}$) the commutation relations (III.4). For ${\mathscr{H}}_{exo}={\mathscr{H}}=-eEx_{2}$ we find,
$\big{(}1-\theta{B^{*}}\big{)}\,\Big{\{}{\mathcal{K}}_{||},{\mathscr{H}}\Big{\}}\propto\Big{\{}x_{2}-\theta\,p_{1}+\hbox{$\textstyle\frac{1}{2}$}eE\,x_{2}^{2},eEx_{2}\Big{\}}\propto-{\theta}\Big{\{}p_{1},x_{2}\Big{\}}=0\,.$
@@
$\left\{\begin{array}[]{lcccl}\mathcal{T}_{1}&=&B^{*}\Big{(}x_{2}-\displaystyle\frac{1}{B^{*}}p_{1}\Big{)}&=&B^{*}Q_{2}\\[8.0pt] \mathcal{T}_{2}&=&-B^{*}\Big{(}x_{1}+\displaystyle\frac{1}{B^{*}}p_{2}-\displaystyle\frac{eE\,}{B^{*}}\,s\Big{)}&=&-B^{*}\Big{(}Q_{1}-\displaystyle\frac{eE\,}{B^{*}}\,s\Big{)}\end{array}\right.$ (V.20)
@@
DJT is the commun projection of both the exo-Galilei and exo-Carroll systems (up to $t\leftrightarrow s$) which thus and inherits both of their symmetries: Galilei boost symmetry is broken, however Carroll boosts along the electric field are symmetries.
@@ VI Motion on the Black Hole horizon
VII Conclusion
The Peierls-Dunne-Jackiw-Trugenberger system (II.1) DunneJP ; DJT , which resurrects Peierls’ historical idea Peierls , can be derived both from the doubly extended (“exotic”) Galilean DHPeierls ; LLGal ; DHJPA ; NCLandau ; ExoRev ; ZHChiral , and from the recently proposed exotic Carroll model Azcarraga ; Ancille0 ; Ancille1 ; Ancille2 ; Marsot21 ; MZH by Hamiltonian reduction FaJa .
The Galilean case is well-known, therefore we focus our attention at the Carroll case. Having suppressed the kinetic term from the Hamiltonian results in a substantially poorer dynamics: the equations for position and for momentum are fully decoupled and the position performs anomalous Hall motion which, for particular values of the parameters, becomes ordinary Hall motion. The extended model allows for an additional, constant “internal” magnetic field represented by our parameter $\kappa_{mag}$ which is added to the external magnetic field $B$ (IV.1) which might play a role in the anomalous Hall effect Kar+ with no external field.
Neither the anomalous and let alone the usual Hall effect can be obtained by an “ordinary” [meaning unextended] Carroll particle. By (IV.7) $\dot{{\bm{x}}}\neq 0$ requires $\theta\neq 0$. The condition (IV.12) requires the constraint (IV.11) which in the massive case $m\neq 0$ can not be satisfied when either of the “exotic” parameters vanishes. When the exotic parameter is turned off, $\theta\to 0$, then eqn (IV.7) behaves as $\dot{x}_{i}\to 0$ confirming the general wisdom that (unextended) Carroll particles do not move.
@@ VII Conclusion
Putting a DJT particle into a planar electromagnatic field breaks Galilei boost symmetry (Proposition V.1). However, owing to the poorer nature of Carroll dynamics, one of the Carroll boosts remains a symmetry for DJT (Proposition V.2) : we get what we called half-Carroll symmetry.
The DJT dynamics actually follows from that of exotic Carroll particles, see Proposition V.5 : half-boost symmetry implies the conservation law (V.17) for ${\mathcal{K}}_{2}$, (Proposition V.17) which, combined with the immobility of $\widetilde{Q}$ in (IV.30), implies in turn no motion along ${\mathbf{E}}$. ${\mathbf{E}}=0$ has arbitrary direction yielding no motion at all.
The “no-motion-property” has recently reemerged, independently, in condensed matter physics in the study of quasiparticles called fractons Pretkofractons ; PretkoCY ; Gromov ; Seiberg ; DoshiG ; Bidussi ; JainJensen ; Grosvenor:2021hkn ; Surowka whose restricted mobility is attributed to their conserved dipole momentum – generated by a dual counterpart of Carroll symmetry MZCH .
Particles with no electric charge but with magnetic momentum and anyonic spin which move on the horizon of a black hole MZH ; Gray behave also analogously to our electrically charged Carroll particles.
Our study shows strong similarity also with vortex dynamics Coslab as it will be explained in a forthcoming paper ZenVort . In conclusion, 450 years after Galilei GalileiDialogo ; Gassendi ; Iglesias the Carrollian context rekindles once again the subtle question of boosts.
@@ Acknowledgements.
We are indebted to X. Bekaert, J. Balog, M. Elbistan, L. Fehér, G. Gibbons, P. Iglesias and F. Ziegler for discussions and for correspondence. PMZ was partially supported by the National Natural Science Foundation of China (Grant No. 11975320).

### v3

@@ V.2 Symmetries of exotic particles
  • Lifting Galilei boosts (V.1) to the evolution space of an exotic Galilean particle,
@@
    the Souriau form (III.5) changes as it does for DJT (V.2),
    $\sigma_{exoG}\to\sigma_{exoG}-eB\epsilon_{ij}b_{i}d{x}_{j}\wedge dt\,,$ (V.13)
    leading to :
@@
    The magnetic field $B\neq 0$ breaks Galilean boost symmetry of an exotic Galilean particle (as seen also directly from in Prop.III.1: the rhs is invariant w.r.t. (V.1)).
@@
$\left\{\begin{array}[]{llcclc}\text{for}&\beta_{1}=0&d\Big{(}\hbox{$\textstyle\frac{1}{2}$}eEx_{2}^{2}\Big{)}\beta_{2}&\text{closed}&{\quad\Rightarrow}&\text{symmetry}\\[8.0pt] \text{for}&\beta_{2}=0&eEx_{1}dx_{2}\,\beta_{1}&\text{not closed}&{\quad\Rightarrow}&{\text{broken}}\end{array}\right.\quad.$ (V.17)
@@
An exotic Carroll particle in the Hall-type electromagnetic field has a half-Carroll symmetry : boosting along the electric field ${\mathbf{E}}=(0,E)$ generates the conserved quantity
${\mathcal{K}}_{||}=(x_{2}-\theta\,p_{1})+\hbox{$\textstyle\frac{1}{2}$}eE\,x_{2}^{2}\,.$ (V.18)
The conservation of ${\mathcal{K}}_{||}$ is readily confirmed by using (after replacing $eB$ by $B^{*}$) the commutation relations (III.4). For ${\mathscr{H}}_{exo}=-eEx_{2}$ we find,
$\big{(}1-\theta{B^{*}}\big{)}\,\Big{\{}{\mathcal{K}}_{||},{\mathscr{H}}_{exo}\Big{\}}\propto\Big{\{}x_{2}-\theta\,p_{1}+\hbox{$\textstyle\frac{1}{2}$}eE\,x_{2}^{2},eEx_{2}\Big{\}}\propto-{\theta}\Big{\{}p_{1},x_{2}\Big{\}}=0\,.$
@@
$\left\{\begin{array}[]{lcccl}\mathcal{T}_{1}&=&B^{*}\Big{(}x_{2}-\displaystyle\frac{1}{B^{*}}p_{1}\Big{)}&=&B^{*}Q_{2}\\[10.0pt] \mathcal{T}_{2}&=&-B^{*}\Big{(}x_{1}+\displaystyle\frac{1}{B^{*}}p_{2}-\displaystyle\frac{eE\,}{B^{*}}\,s\Big{)}&=&-B^{*}\Big{(}Q_{1}-\displaystyle\frac{eE\,}{B^{*}}\,s\Big{)}\end{array}\right.$ (V.21)
@@
DJT is the common projection of both the exotic Galilei and exotic Carroll systems (up to $t\leftrightarrow s$) which thus and inherits both of their symmetries: Galilei boost symmetry is broken, however Carroll boosts along the electric field are symmetries.
@@ VI Motion on the Black Hole horizon
VII Outlook at previous and ongoing research
In this paper, we focus our attention at a constant electromagnetic field in the plane — the usual scenario for the Hall effect. However the theory allows for further physical applications.
Firstly, results (in progress) hint at a Hall-type behavior in a radial electric field, as illustrated by an harmonic DHJPA , or in a newtonian inverse-square force ORPROC ; Peierls2 .
In fact, one can apply the theory to a real physical problem: namely to study a doubly-centrally-extended particle with no magnetic field, $B=0$, submitted to an inverse-square force. In other words, the Kepler problem for an “exotic” particle.
One of the cornerstones of Einstein’s general relativity is the precession of Mercury. The non-commutative Galilean framework provides an alternative description Romero , while in the Carrollian limit the trajectories degenerate to circular ones Precession .
Our results rekindle interest in the long-established theorem of Bertrand Bertrand ; Goldstein which says that the only mechanical systems for which all bounded trajectories are closed are either the harmonic oscillator or the Newtonian inverse-square force. Our forthcoming results Precession hint at that this theorem may require refinement.
Our investigations are rooted in the two-parameter extension of either the Galilei LLGal or of the Carroll Azcarraga ; Ancille0 ; Ancille1 ; Ancille2 ; Marsot21 groups, guaranteed by rigorous mathematical theorems. In the Galilei case, one of the extension parameters is the mass Bargmann . The other one, our $\theta$ here, induces non-commuting coordinates, which has been considered before in the context of the Hall effect Ezawa ; Stone .
For Carroll, the mass is not an extension parameter rather an externally given quantity, however we have again a non-commutativity parameter $\theta$ Azcarraga ; Ancille0 ; Ancille1 ; Ancille2 ; Marsot21 . A most intriguing ingredient is the second parameter $\kappa$, which behaves as a proper magnetic field carried by the particle, which might play a role in explaining the anomalous Hall effect Kar+ . In both cases, the clue is the anomalous velocity relation in (III.2) and in (IV.7), respectively.
Comparison between Carroll and Galilei-invariant theories thus sheds some new light at the two different “non-relativistic” limits: the usual Galilei one, $c\to\infty$, and also the Carrollian one, $c\to 0$ Leblond ; Carrollvs .
We underline, though, that these “exotic” extensions exist only in the plane LLGal ; Azcarraga ; Ancille0 ; Ancille1 ; Ancille2 ; Marsot21 . It is worth to mention, though, that the planar theory has a 3-space-dimensional generalisation, based on the non-Abelian counterpart of the scalar parameter $\theta=\mathop{\rm const.}\nolimits$ promoted to a momentum-dependent vector $\bm{\theta}({\bm{p}})$ Berard1 ; Berard2 . The extended theory can be applied to solid state physics, namely to a Bloch electron in a solid NiuPRL ; NiuRMP . The clue is, once again, an anomalous velocity-momentum relation,
${{\bm{x}}}^{\prime}=\frac{\;{\partial}}{{\partial}{{\bm{p}}}}\epsilon-{{\bm{p}}}^{\prime}\times\bm{\theta}\,,$ (VII.1)
where $\epsilon({\bm{p}})$ is the bound energy NiuPRL ; DHHMSPRL ; DHHMSModPhys ; NiuRMP .
@@ VII Outlook at previous and ongoing research
VIII Conclusion
The Dunne-Jackiw-Trugenberger system (II.1) DunneJP ; DJT which resurrects Peierls’ idea Peierls can be derived from both the doubly extended (“exotic”) Galilean DHPeierls ; LLGal ; DHJPA ; NCLandau ; ExoRev ; ZHChiral , and the recently proposed exotic Carroll model Azcarraga ; Ancille0 ; Ancille1 ; Ancille2 ; Marsot21 ; MZH by Hamiltonian reduction FaJa .
To tackle the position-dependent transformation of Carroll time $s$ in (I.1), we found particularly convenient to work with Souriau’s form (A.5) SSD in Appendix A. However we added also the usual Hamiltonian and Lagrangian formulas for completeness.
The Galilean case is well-known, therefore we focus our attention at the Carroll case. Having suppressed the kinetic term from the Hamiltonian yields a substantially poorer dynamics: the equations for position and for momentum are fully decoupled. The position performs anomalous Hall motion which for particular values of the parameters becomes usual Hall motion.
The extended model allows for an additional, constant “internal” magnetic field represented by our parameter $\kappa_{mag}$ which is added to the external magnetic field $B$ (IV.1). When the non-commutativity parameter does not vanish, $\theta\neq 0$ anomalous Hall motions are found even for $B=0=\kappa_{mag}=0$, cf. (IV.14), see Proposition IV.3. Exotic Carroll particles then could, perhaps, play a role in understanding the anomalous Hall effect which has been observed in the absence of any magnetic field in ferromagnetic material Kar+ ; Ezawa ; Stone .
Neither the anomalous and let alone the usual Hall effect can be obtained by an “ordinary” [meaning unextended] Carroll particle. By (IV.7), $\dot{{\bm{x}}}\neq 0$ requires $\theta\neq 0$. The constraint (IV.11) which, for $m\neq 0$, can not be satisfied when either of the “exotic” parameters vanishes. When the exotic parameter $\theta$ is turned off, then eqn (IV.7) behaves as $\dot{x}_{i}\to 0,$ confirming the general wisdom that (unextended) Carroll particles do not move.
Putting a DJT particle into a planar electromagnetic field breaks Galilei boost symmetry (Proposition V.1). However, owing to the poorer nature of Carroll dynamics, one of the Carroll boosts remains a symmetry for DJT (Proposition V.2) : we get what we called half-Carroll symmetry.
The DJT dynamics actually follows from that of exotic Carroll particles, see Proposition V.5 : half-boost symmetry implies the conservation law (V.18) for ${\mathcal{K}}_{||}$, (Proposition V.18) which implies in turn no motion along ${\mathbf{E}}$.
The “no-motion-property” has recently reemerged, independently, in condensed matter physics namely in the study of quasiparticles called fractons Pretkofractons ; PretkoCY ; Gromov ; Seiberg ; DoshiG ; Bidussi ; JainJensen ; Grosvenor:2021hkn ; Surowka whose restricted mobility is attributed to their conserved dipole momentum, generated by a sort of dual counterpart of Carroll symmetry MZCH .
Particles with no electric charge but with magnetic momentum and anyonic spin which move on the horizon of a black hole MZH ; Gray behave analogously to our electrically charged Carroll particles.
The particle-in-a-field idea can be extended into a curious model LBLL ; Negro ; HMSEnlarged , obtained by “enlarging” the phase space by adding the electromagnetic field (assumed constant), which allows us to restore full Galilei resp. full Carroll boost symmetry ZZZHEnl . Related work on vortex dynamics Coslab ; ORPROC ; ZenVort is in progress. It is somewhat remarkable that 450 years after Galilei GalileiDialogo ; Gassendi ; Iglesias the Carrollian context rekindles the subtle question of boost invariance.
The fundamental question is, though: do Carroll, and in particular “exotic” Carroll particles exist ? The answer should come from experiment. However an argument in its favour might well come from quantum mechanics: massive non-relativistic particles realise the one-parameter central extension of the Galilei group Bargmann . In this paper we propose to adopt the philosophy of Dirac who argued, in the context of magnetic monopoles, “…one would be surprised if Nature had made no use of itDirac .
@@ Acknowledgements.
We are indebted to X. Bekaert, J. Balog, M. Elbistan, L. Fehér, G. Gibbons, P. Iglesias, Z. Silagadze and F. Ziegler for discussions and for correspondence. PMZ is supported by National Key R&D Program of China (Grant No. 2024YFE0109802) and the National Natural Science Foundation of China (Grant No. 12375084).
@@ Appendix A The Noether theorem in Souriau’s framework
The Hall law (II.2) is recovered, for example, from
$\sigma_{DJT}\big{(}({\bm{\xi}}^{\prime},1)\,,\cdot\,\big{)}=e\big{(}B\xi_{i}^{\prime}-\epsilon_{ij}E_{j}\big{)}d\xi_{i}+eE_{i}\xi_{i}^{\prime}ds=0\,,$ (A.2)
where the last term vanishes when the first does.